OpenStudy (babyslapmafro):

Can someone please show me where I went wrong in solving the following problem. Find the local linear approximation of the function f(x)=sqrt(1+x) at x0=0, and use it to approximate sqrt(0.9) and sqrt(1.1)

OpenStudy (anonymous):

your derivative is \(\frac{1}{2\sqrt{x+1}}\)

OpenStudy (babyslapmafro):

Here is what I did... Approximation: sqrt(1) x=0 f(x)=sqrt(1+x)=1 f'(x)=1/[2sqrt(1+x)]=1/2 point (0,1)

OpenStudy (anonymous):

at \(x=0\) your slope is \(\frac{2(x^2+1)-2x(2x)1}{2}\) and your line is \[y=\frac{1}{2}x+1\]

OpenStudy (anonymous):

what on earth???

OpenStudy (babyslapmafro):

y-1=.5(x-0) y=.5(.9-0)+1=1.45 y=.5(1.1-0)+1=1.55

OpenStudy (anonymous):

i meant your slope is \(\frac{1}{2}\)

OpenStudy (anonymous):

so you get \[y=\frac{1}{2}x+1\] now you want to approximate \(\sqrt{.9}\) so if \(x+1=.9\) then \(x=-.1\)

OpenStudy (anonymous):

i see the mistake, you used \(x=.9\) not \(x=-.1\)

OpenStudy (babyslapmafro):

Ok, so I was plugging in the wrong value for x

OpenStudy (anonymous):

yeah, it is \(1+x=.9\) and \(1+x=1.1\) and so \(x=-.1\) and \(x=.1\) respectively

OpenStudy (babyslapmafro):

Ok, then why did they bother giving me the value for x-base 0? It is unnecessary correct?

OpenStudy (babyslapmafro):

I think that value kinda confused me